OpenStudy (anonymous):

The measure of an angle is 78 less than the measure of its complement. what is the measure of the angle?

OpenStudy (anonymous):

90-78 =..

OpenStudy (anonymous):

12

OpenStudy (anonymous):

angle 1 + angle 2 = 90 angle 1 is 78 less than angle 2

OpenStudy (anonymous):

if A is angle 1 and B is angle 2, then A = B-72 And A + B = 90, so (B-72) + B = 90 2B - 72 = 90 2B = 90 + 72 = 162 B = 81 And A is 72 less, or 9. Check work: 9 and 81 are supplementary angles, and 9 is 78 less than 81

OpenStudy (anonymous):

hope that helped...